Find the common factors of the given terms.
6 abc, 24ab^2 , 12a^2 b 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 6abc, 24ab^2and 12a^2b
6abc = 2 × 3 × a × b × c
24 ab^2= 2 × 2 × 2 × 3 × a × b × b
12 a^2b = 2 × 2 × 3 × a × a × b
Common factors of 6 abc, 24ab^2and 12a^2b are 2, 3, a, b and, 2 × 3 × a × b = 6ab

Thanks